# Discovering Coordinated Joint Options via Inter-Agent Relative Dynamics

### v1

@@ 4 Empirical evaluation
Experimental Setup. We evaluated our approach in two multi-agent domains: Level-Based Foraging (papoudakis2021lbf) and Overcooked (ruhdorfer2024overcooked), using their JAX re-implementations from jumanji2024 and Jaxmarl2024, respectively. We focused on two scenarios in each domain: for LBF, these are 15x15-4p-3f and 15x15-4p-5f and for Overcooked, they are Forced Coordination and Counter Circuit. In LBF, we introduced stronger coordination requirements by setting each apple’s level equal to the sum of all agents’ levels, the forced cooperation configuration from (papoudakis2021lbf). Our choice of domains reflects a trade-off between interpretability and feature diversity. LBF enables straightforward visualisation of eigenvectors and relative representations through its $X,Y$-coordinate state space. Overcooked involves richer feature semantics, combining $X,Y$-coordinates, orientations, and a categorical variable for item inventory. While information sharing is implicit in the Overcooked task, in LBF, we allowed agents to always observe the relative distances to their teammates (rather than only when they enter their field of view) and add a flag to their observations indicating when each teammate is in the vicinity of an apple. Our approach and all baselines operated under the same level of observability in our analysis.
@@
To train the $n$-distance encoder and the graph Laplacian eigenvector approximator (ALLO (gomez2024optiondisc)), we used a dataset of 500,000 transitions sampled from a random joint policy in each domain. In LBF, we approximated the first 10 eigenvectors, yielding 20 options, while in Overcooked we approximated 20 eigenvectors, yielding 40 options. Once estimated, we trained joint option policies based on these eigenvectors for one million steps, equivalent to 5% and 10% of the total training time in each task. We employed IQL to train the option policies and incorporated an action, ${\mathcal{A}}^{\prime}_{i}={\mathcal{A}}_{i}\cup\{\bot\}$, as the termination condition (machado2017optiondisc). All agents involved had to all choose this action for an option to be terminated. In addition, we enforced a hard stop after 50 steps. The initiation set was defined as the entire joint state space, allowing options to be started anywhere, provided that the required number of agents, $n_{\mathcal{W}}=N$, was met.
@@
Evaluation against generic baselines. To evaluate H1, we compared the performance of IQL equipped with inter-agent relative options (IQL+IARO) against four option-free baselines: MAPPO (chao2022mappo), IPPO, IQL, and VDN (sunehag@vdn). Following Jaxmarl2024, we left MAPPO out of our analysis for Overcooked. Figure 4 presents IQM scores with 95% confidence intervals (CI) computed across 10 seeds. The addition of joint options (IQL+IARO) led to consistent performance gains over the vanilla IQL, achieving higher percentages of apples eaten per episode in LBF and more successful deliveries per episode in Overcooked; see top row of Figure 4. This improvement was especially visible in Overcooked, where IQL tends to remain stuck in suboptimal solutions. The joint options equipped agents with coordination skills that systematically explored the state space and enabled them to swiftly parse through various coordination patterns in search of better strategies. Additionally, IQL+IARO outperformed the other baselines across the set of experiments, except the Forced Coordination scenario where VDN is known to perform well (Jaxmarl2024). These results highlight the benefits of joint options in overcoming the limitations of independent policy learning in multi-agent tasks by encouraging cooperation through the set of pre-computed coordination behaviours; note that our method does not rely on centralization (MAPPO) or value decomposition (VDN) to support it. However, we did observe a slower convergence at the start of training, which we attribute to known challenges of training with options under global initiation sets (Jong2008TheUO; machado2023optiondisc).
@@
Evaluation against other option frameworks. Next, to evaluate hypothesis H2, we compared the downstream benefits of the options discovered with our framework against those discovered through an existing Kronecker graph product method from KronMAOD2023 (IQL+Kron), and an ablation where discovery was performed directly on raw joint states (IQL+RJS). For IQL+Kron, we closely followed the original implementation, apart from two main adaptations to match our framework: (i) we employed ALLO (gomez2024optiondisc) to approximate both eigenvectors and eigenvalues directly; and (ii) we used a single ROD cycle to estimate multiple eigenvectors. Figure 4 compares the resulting options across both domains and shows that the options discovered by our method better equip teams of agents with the cooperative skills necessary to solve these downstream tasks. We noticed that, particularly in LBF, options discovered via raw joint states and Kronecker products can degrade performance. We believe that this is due to their first non-trivial eigenvectors mainly capturing behaviours that drive agents to the edges of the state space (also see Appendix A.6), which is counterproductive for the apple-picking task.
@@ 5 Related Work
State similarity metrics. Estimating state similarity measures is a fundamental challenge in RL. Even in simple domains, standard distances such as Euclidean, fail to capture true state proximity in the presence of obstacles, as they ignore environmental dynamics. In contrast, bisimulation metrics (ferns2004) measure state similarity through differences in rewards and transition dynamics, and have been widely applied in optimality preserving state aggregation methods (Li2006TowardsAU). However, these approaches struggle in sparse-rewards settings, leaving them susceptible to representation collapse (Kemertas2021; chen2024statechronorepresentationenhancing). Temporal or successor distances define state similarity as the expected number of actions required by a policy to travel between two states (venkattaramanujam2019tempdistance). This class of state distances is invariant to state representations and closely reflects environmental dynamics, making it a popular solution for goal-conditioned RL (hartikainen2020tempdistance; durugkar2021tempdistance; myers2024tempdistance), intrinsic reward composition (bae2024tempdistance; jiang2025tempdistance) and unsupervised skill discovery (park2024tempdistance). Notably, METRA (park2024tempdistance) is particularly relevant to our work, as it aims to learn skills that explore a latent space connected to the ground state space via a temporal metric. Besides performing skill discovery in single-agent settings, a main distinction is that we leverage temporal distances to approximate a latent representation for the eigenoption discovery of joint options that achieve exploration at different time-scales, rather than only at its extremes.
@@
State representations in MARL. State representations in MARL have been mostly used for the aggregation of agents’ local observations into a compact global representation, often through graph neural networks (GNNs). GNNs are invariant to the number of entities, here agent observation embeddings, and can weight the information passed between vertices differently through edge features (jiang2020marlstateabstr; liu2019marlstateabstr; liu2021marlstateabstr; Nayak2023marlstateabstr). utke2025marlstateabstr emphasise the importance of relative information for estimating inter-agent relations, and embed this inductive bias into GNN edge features based on hand-crafted spatial relations. In contrast, our method learns inter-agent relations automatically, without any restrictions on feature semantics.
@@
Temporally extended actions for multi-agent systems. Mahadevan2001masmdp extend semi-Markov decision processes to cooperative multi-agent settings, introducing two execution schemes: synchronous (macro-actions terminate simultaneously) and asynchronous (macro-actions terminate independently). Asynchronous schemes gained recent popularity due to their innate generality (amato2019masmdp; xiao2021masmdp; Xiao2022masmdp), but typically rely on single-agent (local) options that do not express cooperative behaviours. Therefore, coordination occurs only at the option selection level, but not within the option policies themselves. This approach drastically limits the expressiveness of the options used in multi-agent scenarios. To address this, KronMAOD2023 integrates option discovery techniques based on covering options (jinnai2019coveringoptions) that construct joint behaviours via Kronecker products of single-agent transition graphs. However, the resulting joint options mainly synchronise independent behaviours, failing to capture strong inter-agent dependencies or correlations, a limitation noted by the authors. This leaves the problem of discovering strongly coordinated behaviours still open, which is precisely what we aim to address with our method.
@@ A.3 CMI minimisation algorithm
Algorithm 1 follows the framework of dunion2023cmid to minimise the CMI estimator defined in Equation 6.
@@ A.4 Episodic Reward analysis & Additional results
Figure 6 presents an extended analysis of our method using episodic returns as the evaluation measure. Beyond the four scenarios discussed in Section 4, we add one additional setting for each domain: 15x15-3p-5f in Level-Based Foraging (LBF) and Asymmetric Advantages in Overcooked. These scenarios, together with the original four, were selected because the IQL baseline finds it difficult to compute solutions for these scenarios [papoudakis2021lbf, Jaxmarl2024], thereby highlighting the improvements achieved by our approach. Similar to Figure4, we report the IQM scores with 95% CI for 10 seeds.
@@ A.6 Eigenvector comparison
The key difference between eigenvectors derived from the non-relative and relative representations is the lack of responsiveness of the Joint and Kronecker product eigenvectors to different conditioning pairs. This effect is pronounced for the Kronecker product eigenvectors, which remain completely unaffected by teammate positions, further validating the absence of interdependence or coupling in the behaviours captured by this method. This effect arises because both the raw joint state and Kronecker product methods emphasise exploration of the state space rather than inter-agent relations. In contrast, by centring the representation around the Fermat state, the relative representation yields behaviours that adapt to diverse team configurations and produce various patterns of agent alignment. While the scalar representation yields behaviours that align the agents at different distances to each other (on both axes), the multi-dimensional representation enables various in-phase and off-phase behaviours to be discovered with different combinations of features. In addition, as we move deeper into the set of estimated graph Laplacian eigenvectors, they capture progressively shorter time scales. This is a consequence of the orthogonality constraint imposed by estimating multiple eigenvectors per cycle while using graph-drawing–based objectives, ALLO [gomez2024optiondisc].
@@ A.8 Heterogeneous state spaces
In this section, we present our extension of the homogeneous state space framework (Section 3) to heterogeneous state spaces. Due to the need for meaningful state-space alignment, we focus on scenarios in which each agent shares a subset of its state features with one or more teammates. Without such shared features, state synchronisation is not a valid strategy for coordination. Furthermore, we expect each agent to be able to infer its teammates’ types from its observations, either explicitly, when the observations contain type information, or implicitly, for example by enforcing a clear ordering of features or by using fixed positions in the observation vector. We believe these expectations are representative of practical multi-agent collaboration domains, and point to agent modelling techniques [pmlr-v80-rabinowitz18a] as a potential way to relax the latter requirement.
@@
In Section 3, we rely on the temporal successor distance method of myers2024tempdistance for computing pairwise state distances. We now outline how this method can be extended to heterogeneous state spaces through two main changes:
@@
In the original MRN architecture [liu2023temporaldistance_MRN], the distance encoder is computed as a sum of symmetric and asymmetric parts: $d_{\theta}(x,y)=\Delta\!\left(h_{\theta}(x)-h_{\theta}(y)\right)+\left\lVert g_{\theta}(x)-g_{\theta}(y)\right\rVert$, where $\Delta\!(x)=\max_{i=1}^{d}[\max(0,x_{i})]$ and $h_{\theta}$ and $g_{\theta}$ represent the two halves of the outputs of a encoder network parameterised by $\theta$. In our adaptation, we extend this expression to include the agent type as follows. Let $s^{i}_{t}$ and $g^{i}_{t}$ be the source and goal state for agent $i$ at time step $t$, and $l^{i}$ the indexed type of this agent. Then, we modify the above expression for heterogeneous states as: $d_{\theta}(s^{i}_{t},g^{i}_{t},l^{i})=\Delta\!\left(h_{\theta}(s^{i}_{t},l^{i})-h_{\theta}(g^{i}_{t},l^{i})\right)+\left\lVert g_{\theta}(s^{i}_{t},l^{i})-g_{\theta}(g^{i}_{t},l^{i})\right\rVert$. When $s^{i}_{t}$ and $g^{i}_{t}$ correspond to the same state type, $l^{i}$ is redundant. However, as stated in our second proposed change, the padded features of $g^{i}_{t}$, $\{f|f\in{\mathcal{F}}^{*}\wedge f\notin{\mathcal{F}}^{i}\}$, are randomly sampled from the goal states of agents of different types to $i$, resulting in the conditional invariance discussed above.
@@ A.9 ROD cycle discussion
We adopt the original ROD cycle of Machado2017 to approximate eigenoptions, estimating multiple eigenvectors within a single cycle. This design ensures orthogonality among eigenvectors, a property that enables options to function at different time scales. By contrast, other frameworks extract only one eigenvector per cycle, requiring multiple cycles to generate a full set of options, as in covering options [jinnai2019coveringoptions, KronMAOD2023] and covering eigenoptions (CEO) [machado2023optiondisc]. Covering options provide exploration guarantees by leveraging the Fiedler eigenvector to produce policies that connect the farthest points in the state-transition graph, an objective where orthogonality plays little role [jinnai2019coveringoptions]. Our focus, however, is on discovering sets of cooperative behaviours that express diverse alignment patterns. In our experiments with CEO, the most recent framework, we found the resulting eigenvectors to exhibit limited diversity, see Figure 14, reinforcing our decision to return to the original approach.
@@ A.10 Implementation Details & Hyperparameters
$n$-distance training. We followed the publicly available implementation of successor distances from myers2024tempdistance when integrating the CMD-1 architecture into our codebase. The temporal distance encoder $d_{\theta}$ is trained using the symmetrised InfoNCE contrastive loss (without resubstitution) [oord2019infonce], as suggested in the original work. We jointly train this encoder with the Fermat encoder $\phi$ using joint states sampled from a random joint policy and factorised as described in Section 3. Positive pairs for the InfoNCE loss are constructed by pairing current and future states from each agent’s trajectory individually, without cross-agent mixing, while negatives are generated through in-batch shuffling, allowing combinations of states from different agents. To train the Fermat encoder, we incorporate $d_{\theta}$ into the objective in Equation 4 using a stop-gradient operator. When using a multi-dimensional $n$-distance $d_{\theta}^{F}$, we insert a lightweight linear projection to map outputs to a scalar for contrastive loss computation. This projection is discarded after training, restoring the full multi-dimensional relative states. For Fermat encoder training, we instead sum the distance dimensions, since the errors on each dimension are weighted equally for this step.
@@
Optimizer Adam [kingma2015adam] r Adam [kingma2015adam]
@@
ALLO training (Eigenvector approximator). We used the same dataset to train the eigenvector encoder ALLO as for $n$-distance training. We followed the publicly available implementation referenced in gomez2024optiondisc, with similar hyperparameter configurations: two layers of 256 dimensions and barrier coefficient initiation value 2.
@@
Optimizer Adam [kingma2015adam] r Adam [kingma2015adam]
@@
MacDec-POMDP training. We integrated the discovered set of options as additional actions in the original action space of each individual agent. For the backbone IQL implementation, we used an RNN decentralised architecture, trained with the set of hyperparameters presented in Table 3. We used the same list of hyperparameters when training IQL enhanced with options generated from each option discovery framework. To support the training of the decentralised policy over options, we additionally integrated previous SMDP training techniques like training primitive actions on option policy steps, intra-option learning and option interruption [sutton1999optionsdef]. When integrating intra-option learning, we only reused the experiences of the executing joint option to train others when there was an exact match for the actions of each individual agent. For option interruption, we follow the termination-improvement theorem from Sutton1998 and enable an option $w$ to be interrupted if $Q_{i}(H^{i}_{M},{\mathcal{W}})<V(H^{i}_{M})$, for any agent $i$ currently following that option.
@@
Optimizer Adam [kingma2015adam] r Adam [kingma2015adam]

### v2

@@ 4 Empirical evaluation
Experimental Setup. We evaluated our approach in two multi-agent domains: Level-Based Foraging (Papoudakis et al., 2021) and Overcooked (Ruhdorfer et al., 2024), using their JAX re-implementations from 20 and Rutherford et al. (2024), respectively. We focused on two scenarios in each domain: for LBF, these are 15x15-4p-3f and 15x15-4p-5f and for Overcooked, they are Forced Coordination and Counter Circuit. In LBF, we introduced stronger coordination requirements by setting each apple’s level equal to the sum of all agents’ levels, the forced cooperation configuration from (Papoudakis et al., 2021). Our choice of domains reflects a trade-off between interpretability and feature diversity. LBF enables straightforward visualisation of eigenvectors and relative representations through its $X,Y$-coordinate state space. Overcooked involves richer feature semantics, combining $X,Y$-coordinates, orientations, and a categorical variable for item inventory. While information sharing is implicit in the Overcooked task, in LBF, we allowed agents to always observe the relative distances to their teammates (rather than only when they enter their field of view) and add a flag to their observations indicating when each teammate is in the vicinity of an apple. Our approach and all baselines operated under the same level of observability in our analysis.
@@
To train the $n$-distance encoder and the graph Laplacian eigenvector approximator (ALLO (Gomez et al., 2024)), we used a dataset of 500,000 transitions sampled from a random joint policy in each domain. In LBF, we approximated the first 10 eigenvectors, yielding 20 options, while in Overcooked we approximated 20 eigenvectors, yielding 40 options. Once estimated, we trained joint option policies based on these eigenvectors for one million steps, equivalent to 5% and 10% of the total training time in each task. We employed IQL to train the option policies and incorporated an action, ${\mathcal{A}}^{\prime}_{i}={\mathcal{A}}_{i}\cup\{\bot\}$, as the termination condition (Machado et al., 2017b). All agents involved had to all choose this action for an option to be terminated. In addition, we enforced a hard stop after 50 steps. The initiation set was defined as the entire joint state space, allowing options to be started anywhere, provided that the required number of agents, $n_{\mathcal{W}}=N$, was met.
@@
Evaluation against generic baselines. To evaluate H1, we compared the performance of IQL equipped with inter-agent relative options (IQL+IARO) against four option-free baselines: MAPPO (Yu et al., 2022), IPPO, IQL, and VDN (Sunehag et al., 2018). Following Rutherford et al. (2024), we left MAPPO out of our analysis for Overcooked. Figure 4 presents IQM scores with 95% confidence intervals (CI) computed across 10 seeds. The addition of joint options (IQL+IARO) led to consistent performance gains over the vanilla IQL, achieving higher percentages of apples eaten per episode in LBF and more successful deliveries per episode in Overcooked; see top row of Figure 4. This improvement was especially visible in Overcooked, where IQL tends to remain stuck in suboptimal solutions. The joint options equipped agents with coordination skills that systematically explored the state space and enabled them to swiftly parse through various coordination patterns in search of better strategies. Additionally, IQL+IARO outperformed the other baselines across the set of experiments, except the Forced Coordination scenario where VDN is known to perform well (Rutherford et al., 2024). These results highlight the benefits of joint options in overcoming the limitations of independent policy learning in multi-agent tasks by encouraging cooperation through the set of pre-computed coordination behaviours; note that our method does not rely on centralization (MAPPO) or value decomposition (VDN) to support it. However, we did observe a slower convergence at the start of training, which we attribute to known challenges of training with options under global initiation sets (Jong et al., 2008; Machado et al., 2023).
@@
Evaluation against other option frameworks. Next, to evaluate hypothesis H2, we compared the downstream benefits of the options discovered with our framework against those discovered through an existing Kronecker graph product method from Chen et al. (2022) (IQL+Kron), and an ablation where discovery was performed directly on raw joint states (IQL+RJS). For IQL+Kron, we closely followed the original implementation, apart from two main adaptations to match our framework: (i) we employed ALLO (Gomez et al., 2024) to approximate both eigenvectors and eigenvalues directly; and (ii) we used a single ROD cycle to estimate multiple eigenvectors. Figure 4 compares the resulting options across both domains and shows that the options discovered by our method better equip teams of agents with the cooperative skills necessary to solve these downstream tasks. We noticed that, particularly in LBF, options discovered via raw joint states and Kronecker products can degrade performance. We believe that this is due to their first non-trivial eigenvectors mainly capturing behaviours that drive agents to the edges of the state space (also see Appendix A.6), which is counterproductive for the apple-picking task.
@@ 5 Related Work
State similarity metrics. Estimating state similarity measures is a fundamental challenge in RL. Even in simple domains, standard distances such as Euclidean, fail to capture true state proximity in the presence of obstacles, as they ignore environmental dynamics. In contrast, bisimulation metrics (Ferns et al., 2004) measure state similarity through differences in rewards and transition dynamics, and have been widely applied in optimality preserving state aggregation methods (Li et al., 2006). However, these approaches struggle in sparse-rewards settings, leaving them susceptible to representation collapse (Kemertas and Aumentado-Armstrong, 2021; Chen et al., 2024). Temporal or successor distances define state similarity as the expected number of actions required by a policy to travel between two states (Venkattaramanujam et al., 2019). This class of state distances is invariant to state representations and closely reflects environmental dynamics, making it a popular solution for goal-conditioned RL (Hartikainen et al., 2020; Durugkar et al., 2021; Myers et al., 2024), intrinsic reward composition (Bae et al., 2024; Jiang et al., 2025) and unsupervised skill discovery (Park et al., 2024). Notably, METRA (Park et al., 2024) is particularly relevant to our work, as it aims to learn skills that explore a latent space connected to the ground state space via a temporal metric. Besides performing skill discovery in single-agent settings, a main distinction is that we leverage temporal distances to approximate a latent representation for the eigenoption discovery of joint options that achieve exploration at different time-scales, rather than only at its extremes.
@@
State representations in MARL. State representations in MARL have been mostly used for the aggregation of agents’ local observations into a compact global representation, often through graph neural networks (GNNs). GNNs are invariant to the number of entities, here agent observation embeddings, and can weight the information passed between vertices differently through edge features (Jiang et al., 2020; Liu et al., 2019, 2021; Nayak et al., 2023). Utke et al. (2025) emphasise the importance of relative information for estimating inter-agent relations, and embed this inductive bias into GNN edge features based on hand-crafted spatial relations. In contrast, our method learns inter-agent relations automatically, without any restrictions on feature semantics.
@@
Temporally extended actions for multi-agent systems. Makar et al. (2001) extend semi-Markov decision processes to cooperative multi-agent settings, introducing two execution schemes: synchronous (macro-actions terminate simultaneously) and asynchronous (macro-actions terminate independently). Asynchronous schemes gained recent popularity due to their innate generality (Amato et al., 2019; Xiao et al., 2021, 2022), but typically rely on single-agent (local) options that do not express cooperative behaviours. Therefore, coordination occurs only at the option selection level, but not within the option policies themselves. This approach drastically limits the expressiveness of the options used in multi-agent scenarios. To address this, Chen et al. (2022) integrates option discovery techniques based on covering options (Jinnai et al., 2019a) that construct joint behaviours via Kronecker products of single-agent transition graphs. However, the resulting joint options mainly synchronise independent behaviours, failing to capture strong inter-agent dependencies or correlations, a limitation noted by the authors. This leaves the problem of discovering strongly coordinated behaviours still open, which is precisely what we aim to address with our method.
@@ A.3 CMI minimisation algorithm
Algorithm 1 follows the framework of Dunion et al. [2023] to minimise the CMI estimator defined in Equation 6.
@@ A.4 Episodic Reward analysis & Additional results
Figure 6 presents an extended analysis of our method using episodic returns as the evaluation measure. Beyond the four scenarios discussed in Section 4, we add one additional setting for each domain: 15x15-3p-5f in Level-Based Foraging (LBF) and Asymmetric Advantages in Overcooked. These scenarios, together with the original four, were selected because the IQL baseline finds it difficult to compute solutions for these scenarios [Papoudakis et al., 2021, Rutherford et al., 2024], thereby highlighting the improvements achieved by our approach. Similar to Figure4, we report the IQM scores with 95% CI for 10 seeds.
@@ A.6 Eigenvector comparison
The key difference between eigenvectors derived from the non-relative and relative representations is the lack of responsiveness of the Joint and Kronecker product eigenvectors to different conditioning pairs. This effect is pronounced for the Kronecker product eigenvectors, which remain completely unaffected by teammate positions, further validating the absence of interdependence or coupling in the behaviours captured by this method. This effect arises because both the raw joint state and Kronecker product methods emphasise exploration of the state space rather than inter-agent relations. In contrast, by centring the representation around the Fermat state, the relative representation yields behaviours that adapt to diverse team configurations and produce various patterns of agent alignment. While the scalar representation yields behaviours that align the agents at different distances to each other (on both axes), the multi-dimensional representation enables various in-phase and off-phase behaviours to be discovered with different combinations of features. In addition, as we move deeper into the set of estimated graph Laplacian eigenvectors, they capture progressively shorter time scales. This is a consequence of the orthogonality constraint imposed by estimating multiple eigenvectors per cycle while using graph-drawing–based objectives, ALLO [Gomez et al., 2024].
@@ A.8 Heterogeneous state spaces
In this section, we present our extension of the homogeneous state space framework (Section 3) to heterogeneous state spaces. Due to the need for meaningful state-space alignment, we focus on scenarios in which each agent shares a subset of its state features with one or more teammates. Without such shared features, state synchronisation is not a valid strategy for coordination. Furthermore, we expect each agent to be able to infer its teammates’ types from its observations, either explicitly, when the observations contain type information, or implicitly, for example by enforcing a clear ordering of features or by using fixed positions in the observation vector. We believe these expectations are representative of practical multi-agent collaboration domains, and point to agent modelling techniques [Rabinowitz et al., 2018] as a potential way to relax the latter requirement.
@@
In Section 3, we rely on the temporal successor distance method of Myers et al. [2024] for computing pairwise state distances. We now outline how this method can be extended to heterogeneous state spaces through two main changes:
@@
In the original MRN architecture [Liu et al., 2023], the distance encoder is computed as a sum of symmetric and asymmetric parts: $d_{\theta}(x,y)=\Delta\!\left(h_{\theta}(x)-h_{\theta}(y)\right)+\left\lVert g_{\theta}(x)-g_{\theta}(y)\right\rVert$, where $\Delta\!(x)=\max_{i=1}^{d}[\max(0,x_{i})]$ and $h_{\theta}$ and $g_{\theta}$ represent the two halves of the outputs of a encoder network parameterised by $\theta$. In our adaptation, we extend this expression to include the agent type as follows. Let $s^{i}_{t}$ and $g^{i}_{t}$ be the source and goal state for agent $i$ at time step $t$, and $l^{i}$ the indexed type of this agent. Then, we modify the above expression for heterogeneous states as: $d_{\theta}(s^{i}_{t},g^{i}_{t},l^{i})=\Delta\!\left(h_{\theta}(s^{i}_{t},l^{i})-h_{\theta}(g^{i}_{t},l^{i})\right)+\left\lVert g_{\theta}(s^{i}_{t},l^{i})-g_{\theta}(g^{i}_{t},l^{i})\right\rVert$. When $s^{i}_{t}$ and $g^{i}_{t}$ correspond to the same state type, $l^{i}$ is redundant. However, as stated in our second proposed change, the padded features of $g^{i}_{t}$, $\{f|f\in{\mathcal{F}}^{*}\wedge f\notin{\mathcal{F}}^{i}\}$, are randomly sampled from the goal states of agents of different types to $i$, resulting in the conditional invariance discussed above.
@@ A.9 ROD cycle discussion
We adopt the original ROD cycle of Machado et al. [2017a] to approximate eigenoptions, estimating multiple eigenvectors within a single cycle. This design ensures orthogonality among eigenvectors, a property that enables options to function at different time scales. By contrast, other frameworks extract only one eigenvector per cycle, requiring multiple cycles to generate a full set of options, as in covering options [Jinnai et al., 2019a, Chen et al., 2022] and covering eigenoptions (CEO) [Machado et al., 2023]. Covering options provide exploration guarantees by leveraging the Fiedler eigenvector to produce policies that connect the farthest points in the state-transition graph, an objective where orthogonality plays little role [Jinnai et al., 2019a]. Our focus, however, is on discovering sets of cooperative behaviours that express diverse alignment patterns. In our experiments with CEO, the most recent framework, we found the resulting eigenvectors to exhibit limited diversity, see Figure 14, reinforcing our decision to return to the original approach.
@@ A.10 Implementation Details & Hyperparameters
$n$-distance training. We followed the publicly available implementation of successor distances from Myers et al. [2024] when integrating the CMD-1 architecture into our codebase. The temporal distance encoder $d_{\theta}$ is trained using the symmetrised InfoNCE contrastive loss (without resubstitution) [van den Oord et al., 2019], as suggested in the original work. We jointly train this encoder with the Fermat encoder $\phi$ using joint states sampled from a random joint policy and factorised as described in Section 3. Positive pairs for the InfoNCE loss are constructed by pairing current and future states from each agent’s trajectory individually, without cross-agent mixing, while negatives are generated through in-batch shuffling, allowing combinations of states from different agents. To train the Fermat encoder, we incorporate $d_{\theta}$ into the objective in Equation 4 using a stop-gradient operator. When using a multi-dimensional $n$-distance $d_{\theta}^{F}$, we insert a lightweight linear projection to map outputs to a scalar for contrastive loss computation. This projection is discarded after training, restoring the full multi-dimensional relative states. For Fermat encoder training, we instead sum the distance dimensions, since the errors on each dimension are weighted equally for this step.
@@
Optimizer Adam [Kingma and Ba, 2017] r Adam [Kingma and Ba, 2017]
@@
ALLO training (Eigenvector approximator). We used the same dataset to train the eigenvector encoder ALLO as for $n$-distance training. We followed the publicly available implementation referenced in Gomez et al. [2024], with similar hyperparameter configurations: two layers of 256 dimensions and barrier coefficient initiation value 2.
@@
Optimizer Adam [Kingma and Ba, 2017] r Adam [Kingma and Ba, 2017]
@@
MacDec-POMDP training. We integrated the discovered set of options as additional actions in the original action space of each individual agent. For the backbone IQL implementation, we used an RNN decentralised architecture, trained with the set of hyperparameters presented in Table 3. We used the same list of hyperparameters when training IQL enhanced with options generated from each option discovery framework. To support the training of the decentralised policy over options, we additionally integrated previous SMDP training techniques like training primitive actions on option policy steps, intra-option learning and option interruption [Sutton et al., 1999]. When integrating intra-option learning, we only reused the experiences of the executing joint option to train others when there was an exact match for the actions of each individual agent. For option interruption, we follow the termination-improvement theorem from Sutton and Barto [1998] and enable an option $w$ to be interrupted if $Q_{i}(H^{i}_{M},{\mathcal{W}})<V(H^{i}_{M})$, for any agent $i$ currently following that option.
@@
Optimizer Adam [Kingma and Ba, 2017] r Adam [Kingma and Ba, 2017]